If the diameter of a cylindrical can is increased by 30 percent, by approximating what percentage should the height be increased to triple the volume of the can ?
Answer by stanbon(75887)

If the diameter of a cylindrical can is increased by 30 percent, by approximating what percentage should the height be increased to triple the volume of the can ?
-----------
Before Increase:
V = (pi*r^2)h
---
After Increase:
3V = (pi*r^2)H
-------------------
H = 3V/(pi*r^2)
H = [3(pi*r^2)h]/(pi*r^2)
H = 3h
H/h = 300%
======================
Cheers,
Stan H.